# Unforgettable Lessons from Forgettable Images: Intra-Class Memorability Matters in Computer Vision Tasks

### v1

@@ Abstract
We introduce intra-class memorability, where certain images within the same class are more memorable than others despite shared category characteristics. To investigate what features make one object instance more memorable than others, we design and conduct human behavior experiments, where participants are shown a series of images one at a time, and they must identify when the current item matches the item presented a few steps back in the sequence. To quantify memorability, we propose the Intra-Class Memorability score (ICMscore), a novel metric that incorporates the temporal intervals between repeated image presentations into its calculation. Furthermore, we curate the Intra-Class Memorability Dataset (ICMD), comprising over 2,500 images across five object classes with their ICMscores derived from 1,000 participants’ responses in the behavioral experiment. Subsequently, we demonstrate the usefulness of ICMD by training AI models on this dataset for various downstream tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Our findings show that models trained on ICMD outperform existing memorability predictors, capturing visual features essential for intra-class memorability. Surprisingly, high-ICMscore images impair AI performance in image recognition and continual learning tasks, while low-ICMscore images improve outcomes in these tasks. Additionally, we fine-tune a state-of-the-art image diffusion model on ICMD image pairs with and without masked semantic objects. The diffusion model can successfully manipulate image elements to enhance or reduce memorability. Our contributions open new pathways in understanding intra-class memorability by scrutinizing fine-grained visual features that result in the least and most memorable images and laying the groundwork for real-world applications in cognitive science and computer vision. All code, data, and models will be made publicly available.
Figure 1: An object instance can be more memorable than another within the same class. Intra-class memorability refers to the ability of instances within the same category to be remembered. First, we present example images with low and high memorability (first and second rows) across five object classes (five columns). For example, in Column 1, a plain white teapot is less memorable than a rainbow-colored teapot. Images are identified as high or low memorability based on results from human behavior experiments. We then use these high-memorability and low-memorability images to train and test state-of-the-art AI models across four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Each task is highlighted in dashed bubbles with distinct colors.
@@ 1 Introduction
As we experience the world, our memory selectively retains certain objects while overlooking others [35, 18, 32] (Fig. 1). For instance, among numerous flowers we encounter, a pale white daisy might fade from memory, while a striking violet osteospermum with intricate shapes remains vividly etched (Column 3 of Fig. 1). What factors make some flowers more memorable than others? To answer this question, we first introduce intra-class memorability which refers to the ability of instances within the same object class to be remembered.
To delve into why certain items are more memorable than others, despite belonging to the same overall category, we conduct human behavior experiments using a continuous recognition task. In this task, participants are presented with a sequence of new and repeated images from the same class with delayed repetition and they are asked to identify those repeats. The proposed paradigm is novel compared to existing research, where continuous recognition tasks typically involve sequential presentations of images from different classes [7, 26, 13, 15, 30, 21].
To quantitatively measure the memorability of individual instances, we introduce the Intra-Class Memorability score (ICMscore). Unlike prior evaluation metrics [15, 30, 24, 1], our ICMscore incorporates average hit and miss rates for an item’s repetitions, weighted by the time intervals between its initial and repeated presentations. Missing a repeat after a short interval indicates low memorability (LM), while remembering a repeat after a long interval signifies high memorability (HM).
@@
Our experimental results show that a memorability prediction model trained on ICMD captures representative features diagnostic for intra-class memorability. Moreover, surprisingly, while highly memorable images seem more salient and thus easier to remember, we found that these images actually impair AI performance in image recognition and continual learning tasks. Finally, we use images from ICMD with and without masked semantic objects to fine-tune the state-of-the-art image diffusion model [6] using LoRA [22]. The diffusion model can manipulate elements in the scene to control image memorability. We highlight our key contributions:
@@
  • We challenge existing paradigms for collecting memorability data on naturalistic images by introducing a new human behavioral experiment on intra-class memorability. Additionally, we propose a novel memorability metric, the ICMscore, which accounts for the time intervals before a repeated image is missed or identified in continuous image recognition.
  • We collect ICMscore from the responses of 1000 participants on 2500 images spanning across 5 object classes and contribute the first dataset on intra-class memorability, named as Intra-Class Memorability dataset (ICMD). To demonstrate the usefulness of our dataset, we introduce four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing.
@@ 2 Related Work
Research in cognitive psychology indicates that certain visual and contextual features—such as prominent visual characteristics, emotional valence, and semantic richness—substantially increase the likelihood of remembering specific objects over others [44]. This implies that memorability is a predictable and universal property, rather than purely subjective [29, 8, 23]. Based on this, Isola et al. [23] define “memorability” as the probability that a viewer will remember an image after a single viewing. Subsequent studies have investigated how intrinsic and extrinsic factors influence memorability. Extrinsic factors refer to an image’s categorical information, while intrinsic factors include other visual features beyond these categorical information [26, 13, 21]. However, none of these studies have effectively disentangled the effects of extrinsic and intrinsic features on memorability.
@@
To address this problem, subsequent works [7, 15, 30, 34, 5] focus on specific super-categories, such as animals or indoor and outdoor scenes. In addition, some research has extended the studies on static images to videos, exploring memorability differences across video super-categories [9, 38, 10]. Although these studies limit categories to reduce the impact of extrinsic factors on memorability, they still mix images from different sub-categories, making it difficult to conduct controlled studies that solely examine the effect of intrinsic features on memorability. To address this research gap, we introduce the concept of intra-class memorability. To investigate it, we propose a novel human behavior experimental paradigm to measure memorability differences among images within the same class.
@@
Several memorability datasets have been collected through human behavior experiments [23, 7, 13, 15, 30], with LaMem [26] as one of the largest for deep learning-based image memorability prediction. In these studies, participants typically view a sequence of images, identifying any repeated target images. Memorability scores are often calculated based on the proportion of participants who correctly identify these repeats [23, 7, 13]. Some studies adjust scores to account for false alarms, where images are mistakenly identified as repeats [4, 15, 30]. The time interval between target presentations also influences memorability [26], yet random sampling of intervals can lead to unfair comparisons due to inconsistent controls. In our experiment, we standardize time intervals and introduce a novel memorability score, named as Intra-Class Memorability score (ICMscore). It accounts for both hit rates and miss rates of repeated images across a predefined set of time intervals. Moreover, we contribute a new dataset, Intra-Class Memorability dataset (ICMD). It consists of 2500 images spanning five object classes with their corresponding ICMscores based on responses from over 1000 participants.
@@
Several studies have demonstrated that HM images can positively influence human cognitive states. For example, GANalyze [16] introduces a generative model that modifies image attributes to affect human emotional and cognitive responses. Peng et al. [41] show that memorable images boost user engagement in image retrieval systems and improve recall accuracy in storytelling and educational contexts. Similarly, Pataranutaporn et al. [40] explore how AI-modified images and videos uniquely shape human memory. Thus, predicting memorability for a given image has become increasingly important [31, 49, 17]. However, existing models for predicting memorability rely on both extrinsic and intrinsic features, making them unable to capture fine-grained memorability differences within the same class of images. In contrast, our experiments demonstrate that our model, trained on the ICMD dataset, outperforms existing predictors in accurately predicting intra-class memorability.
@@
Despite early efforts to explore the role of memorability in enhancing human cognition, its impact on computer vision tasks remains largely unexplored. Specifically, the connection between image features relevant to object recognition and those that explain memorability is still unclear. Understanding this relationship could have significant implications for fields like continual learning. Using our ICMD dataset, we present preliminary evidence that memorability affects AI performance in tasks such as image recognition and continual learning. To edit images with controls on memorability, we also build upon the existing generative models [6, 22] and introduce an image diffusion model, capable of manipulating image elements to enhance or reduce memorability.
@@ 3.1 Human Behavior Experiments
Different from the existing memorability datasets [7, 26, 13, 21, 15, 30], we focus on collecting a new Intra-Class Memorability dataset (ICMD), which features naturalistic images from five object classes and their associated human memorability responses. Fig. 2 presents the schematic of the human psychophysics experiments in the continuous object recognition tasks. In each trial, participants view a sequence of images, with both target and foil images belonging to the same category. Each image is followed by a blank screen to allow ample response time. Participants press the space bar when recognizing a previously seen image, with reaction times recorded during the 1,200 ms image presentation and the subsequent 1,600 ms blank screen. Before the start of an experiment, participants complete a guided demonstration to ensure task comprehension. During the experiment, immediate feedback follows each response: a green tick for correct answers and a red cross for incorrect ones. The image sequence is randomly shuffled, with target images reappearing after intervals of $t\in[8,16,24,32]$ images, uniformly distributed across these intervals. Each trial includes 100 unique target images, with foil images filling the sequence when targets are not shown. To maintain data quality, each target image and its repeat are shown only once to one participant. Sec. S1 and Fig. S1 highlight differences in the experiment schematic between our study and previous works.
We select 2,500 images from ImageNet [11] as target images for the experiment, evenly distributed across five categories: dogs, cars, flowers, jellyfish, and teapots. The experiment is conducted on Amazon Mechanical Turk (MTurk) [51] with 3,263 participants. To ensure data quality, participants are disqualified if they incorrectly identify more than three foil images as target images in a row or fail to respond within 60 seconds. After filtering the data from disqualified participants, we analyzed responses from the remaining 1,000 qualified participants, resulting in 2,000 trials and a total of 171,568 key presses, with an average of 86 key presses per trial. The average false positive rate of the key presses is 14.67%. All experiments are conducted with informed consent and follow protocols approved by our institution’s Institutional Review Board.
@@ 3.2 Intra-Class Memorability Score (ICMscore)
To assess intra-class memorability, we introduce the Intra-Class Memorability Score (ICMscore), which measures the memorability of an image independently of its categorical information. Previous research typically computes memorability scores using hit rates and miss rates of memory recall [4, 15, 30]. However, the interval $t$ between image repetitions significantly affects memorability [26]. As shown in Fig. S2 and Sec. S2, memory accuracy declines with longer intervals due to memory decays and increasing cognitive loads over time [52, 48]. Therefore, we incorporate the lengths of these time intervals into the ICMscore calculation. Next, we describe how we compute ICMscore for a target image $I$.
@@
Since human memory retention diminishes over longer intervals, a longer interval $t$ makes correct identification of seen images more challenging. Therefore, when a participant correctly identifies an image after a longer interval, we assign it a higher memorability score $s$ to reflect the increased difficulty associated with the larger $t$. For correct responses ($x_{i}=1$), the memorability score $s_{i}$ is computed using a linearly increasing function: $s_{i}=x_{i}\frac{t_{i}}{T}$, where $t_{i}$ is the interval between two presentations of a target image, and $T=32$ is the maximum interval.
Conversely, as longer time intervals are more prone to incorrect identifications of repeated target images, errors during shorter intervals are more indicative of the image’s memorability. For incorrect responses ($x_{i}=-1$), shorter intervals incur a larger penalty, while longer intervals result in a smaller penalty. The memorability score for incorrect responses is given by $s_{i}=x_{i}\frac{T-t_{i}}{T}$. The final ICMscore for each target image $I$ is calculated as the average of its memorability scores $s_{i}$ over all $n$ participants: ICMscore(I)$=\frac{\sum_{i=1}^{n}s_{i}}{n}$.
@@ 3.3 Dataset Analysis
Given the target images and their corresponding ICMscores, we now analyze how visual features of target images influence ICMscores. To investigate this relationship, we use a ResNet [20] model pre-trained on ImageNet [11] to extract feature representations for each image. We define the mean of all image features belonging to the same category as the centroid. Next, we compute the Pearson correlation [45] between the L2 distances from its centroid to individual target image features and their ICMscores.
Figure 3: Correlation analysis between the feature distances of individual images to their categorical centroids and their corresponding ICMscore for the car and flower categories. Each “Car” and “Flower” image in ICMD is represented as a scatter point, with its corresponding ICMscore shown on the y-axis and the L2 distance from their centroids on the x-axis. The dashed lines, derived from linear regression, indicate a positive correlation between feature distance and CIMscore. As the L2 distance from the centroid increases, the ICMscore generally increases. The Pearson correlation coefficients are $r=0.46$ for “Car” and $r=0.42$ for “Flower” ($p<0.05$). See Sec. S3 for the correlation analysis on other object categories.
Our results reveal a positive correlation between an image’s distance from its categorical centroid and its ICMscore across all object classes, as shown in Fig. 3 for the car and flower classes and Fig. S3 for the other object classes. The further away the features of an image are from its categorical centroid, the more memorable the image becomes. For example, a visually striking flower with distinct colors or intricate petal patterns stands out from typical flowers, making it more likely to be remembered (Column 1 in Fig. 1). Here, we explored the connections between memorability and feature distance within categories. Insights gained from this intra-class feature analysis motivate us to design the experiments that investigate the role of memorability in the following four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing.

### v2

@@ Abstract
We introduce intra-class memorability, where certain images within the same class are more memorable than others despite shared category characteristics. To investigate what features make one object instance more memorable than others, we design and conduct human behavior experiments, where participants are shown a series of images, and they must identify when the current image matches the image presented a few steps back in the sequence. To quantify memorability, we propose the Intra-Class Memorability score (ICMscore), a novel metric that incorporates the temporal intervals between repeated image presentations into its calculation. Furthermore, we curate the Intra-Class Memorability Dataset (ICMD), comprising over 2,500 images across five object classes with their ICMscores derived from 1,000 participants’ responses. Subsequently, we demonstrate the usefulness of ICMD by training AI models on this dataset for various downstream tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Surprisingly, high-ICMscore images impair AI performance in image recognition and continual learning tasks, while low-ICMscore images improve outcomes in these tasks. Additionally, we fine-tune a state-of-the-art image diffusion model on ICMD image pairs with and without masked semantic objects. The diffusion model can successfully manipulate image elements to enhance or reduce memorability. Our contributions open new pathways in understanding intra-class memorability by scrutinizing fine-grained visual features that result in the least and most memorable images and laying the groundwork for real-world applications in cognitive science and computer vision. All code, data, and models will be made publicly available.
Cognitive Science, Memorability
@@ 1 Introduction
Figure 1: An object instance can be more memorable than another within the same class. Intra-class memorability refers to the ability of instances within the same category to be remembered. First, we present example images with low and high memorability (first and second rows) for the exemplar flower class. A white pale daisy (Row 1) is less memorable than a striking violet osteospermum (Row 2). Images are identified as high or low memorability based on results from human behavior experiments. We then use these high-memorability and low-memorability images to train and test state-of-the-art AI models across four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Each task is highlighted in dashed bubbles with distinct colors.
As we experience the world, our memory selectively retains certain objects while overlooking others (McDermott & Roediger, 2018; Halamish & Stern, 2022; Leshikar et al., 2012). For instance, among numerous flowers we encounter, a pale white daisy might fade from memory, while a striking violet osteospermum with intricate shapes remains vividly etched (Fig. 1). What factors make some flowers more memorable than others? To answer this question, we first introduce intra-class memorability which refers to the ability of instances within the same object class to be remembered.
To delve into why certain items are more memorable than others, despite belonging to the same overall category, we conduct human behavior experiments using a continuous recognition task. In this task, participants are presented with a sequence of new and repeated images from the same class with delayed repetition and they are asked to identify those repeats. The proposed paradigm is novel compared to existing research, where continuous recognition tasks typically involve sequential presentations of images from different classes (Bylinskii et al., 2015; Khosla et al., 2015; Dubey et al., 2015; Goetschalckx & Wagemans, 2019; Kramer et al., 2023; Hovhannisyan et al., 2021).
To measure the memorability of individual instances, we introduce the Intra-Class Memorability score (ICMscore). Unlike prior evaluation metrics (Goetschalckx & Wagemans, 2019; Kramer et al., 2023; Isola et al., 2014; Akagunduz et al., 2019), our ICMscore incorporates average hit and miss rates for an item’s repetitions, weighted by the time intervals between its initial and repeated presentations. Missing a repeat after a short interval indicates low memorability (LM), while remembering a repeat after a long interval signifies high memorability (HM).
@@
Our experimental results show that a memorability prediction model trained on ICMD captures representative features diagnostic for intra-class memorability. Moreover, surprisingly, while highly memorable images seem more salient and thus easier to remember, we found that these images actually impair AI performance in image recognition and continual learning tasks. Finally, we use images from ICMD with and without masked semantic objects to fine-tune the state-of-the-art image diffusion model (Brooks et al., 2022) using LoRA (Hu et al., 2021). The diffusion model can edit image elements to control memorability.
We highlight our key contributions:
@@
  • We challenge existing paradigms for collecting memorability data on naturalistic images by introducing a new human behavioral experiment on intra-class memorability. In addition, we propose a novel memorability metric, the ICMscore, which accounts for the time intervals before a repeated image is identified in continuous image recognition.
  • We collect ICMscore from the responses of 1000 participants on 2500 images spanning across 5 object classes and contribute the first dataset on intra-class memorability, named as Intra-Class Memorability dataset (ICMD). To demonstrate the usefulness of our dataset, we introduce four computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing.
@@ 2 Related Work
Research in cognitive psychology indicates that certain visual and contextual features—such as prominent visual characteristics, emotional valence, and semantic richness—substantially increase the likelihood of remembering specific objects over others (Schmidt, 1991). This implies that memorability is a predictable and universal property, rather than purely subjective (Konkle et al., 2010; Bylinskii et al., 2022; Isola et al., 2011). Based on this, Isola et al. (Isola et al., 2011) define “memorability” as the probability that a viewer will remember an image after a single viewing. Subsequent studies have investigated how intrinsic and extrinsic factors influence memorability. Extrinsic factors refer to an image’s categorical information, while intrinsic factors include other visual features beyond these categorical information (Khosla et al., 2015; Dubey et al., 2015; Hovhannisyan et al., 2021). However, none of these studies have effectively disentangled the effects of extrinsic and intrinsic features on memorability.
@@
To address this problem, subsequent works (Bylinskii et al., 2015; Goetschalckx & Wagemans, 2019; Kramer et al., 2023; Lu et al., 2020; Borkin et al., 2013) focus on specific super-categories, such as animals or indoor and outdoor scenes. In addition, some research has extended the studies on static images to videos, exploring memorability differences across video super-categories (Cohendet et al., 2018a; Newman et al., 2020; Cohendet et al., 2018b). Although these studies limit categories to reduce the impact of extrinsic factors on memorability, they still mix images from different sub-categories, making it difficult to conduct controlled studies that solely examine the effect of intrinsic features on memorability. To address this research gap, we introduce the concept of intra-class memorability. To investigate it, we propose a novel human behavior experimental paradigm to measure memorability differences among images within the same class.
@@
Several memorability datasets have been collected through human behavior experiments (Isola et al., 2011; Bylinskii et al., 2015; Dubey et al., 2015; Goetschalckx & Wagemans, 2019; Kramer et al., 2023), with LaMem (Khosla et al., 2015) as one of the largest for deep learning-based image memorability prediction. In these studies, participants typically view a sequence of images, identifying any repeated target images. Memorability scores are often calculated based on the proportion of participants who correctly identify these repeats (Isola et al., 2011; Bylinskii et al., 2015; Dubey et al., 2015). Some studies adjust scores to account for false alarms, where images are mistakenly identified as repeats (Bainbridge & Rissman, 2018; Goetschalckx & Wagemans, 2019; Kramer et al., 2023). The time interval between target presentations also influences memorability (Khosla et al., 2015), yet random sampling of intervals can lead to unfair comparisons due to inconsistent controls. In our experiment, we standardize time intervals and introduce a novel memorability score, named as Intra-Class Memorability score (ICMscore). It accounts for both hit rates and miss rates of repeated images across a predefined set of time intervals. Moreover, we contribute a new dataset, Intra-Class Memorability dataset (ICMD). It consists of 2500 images spanning five object classes with their corresponding ICMscores based on responses from over 1000 participants.
@@
Several studies have demonstrated that HM images can positively influence human cognitive states. For example, GANalyze (Goetschalckx et al., 2019) introduces a generative model that modifies image attributes to affect human emotional and cognitive responses. Peng et al. (Peng & Bainbridge, 2024) show that memorable images boost user engagement in image retrieval systems and improve recall accuracy in storytelling and educational contexts. Similarly, (Pataranutaporn et al., 2024) and (Siarohin et al., 2019) explore how AI-modified images and videos uniquely shape human memory. Thus, predicting memorability for a given image has become increasingly important (Lahrache & El Ouazzani, 2022; Squalli-Houssaini et al., 2018; Hagen & Espeseth, 2023). However, existing models for predicting memorability rely on both extrinsic and intrinsic features, making them unable to capture fine-grained memorability differences within the same class of images. In contrast, our experiments demonstrate that our model, trained on the ICMD dataset, outperforms existing predictors in accurately predicting intra-class memorability.
@@
Despite early efforts to explore the role of memorability in enhancing human cognition, its impact on computer vision tasks remains largely unexplored. Specifically, the connection between image features relevant to object recognition and those that explain memorability is still unclear. Understanding this relationship could have significant implications for fields like continual learning. Using our ICMD dataset, we present preliminary evidence that memorability affects AI performance in tasks such as image recognition and continual learning. To edit images with controls on memorability, we also build upon the existing generative models (Brooks et al., 2022; Hu et al., 2021) and introduce an image diffusion model, capable of manipulating image elements to enhance or reduce memorability.
@@ 3.1 Human Behavior Experiments
Different from the existing memorability datasets (Bylinskii et al., 2015; Khosla et al., 2015; Dubey et al., 2015; Hovhannisyan et al., 2021; Goetschalckx & Wagemans, 2019; Kramer et al., 2023), we focus on collecting a new Intra-Class Memorability dataset (ICMD), which features naturalistic images from five object classes and their associated human memorability responses. Fig. 2 presents the schematic of the human psychophysics experiments in the continuous object recognition tasks. In each trial, participants view a sequence of images, with both target and foil images belonging to the same category. Each image is followed by a blank screen to allow ample response time. Participants press the space bar when recognizing a previously seen image, with reaction times recorded during the 1,200 ms image presentation and the subsequent 1,600 ms blank screen. Before the start of an experiment, participants complete a guided demonstration to ensure task comprehension. During the experiment, immediate feedback follows each response: a green tick for correct answers and a red cross for incorrect ones. The image sequence is randomly shuffled, with target images reappearing after intervals of $t\in[8,16,24,32]$ images, uniformly distributed across these intervals. We determine these interval ranges through a pilot study exploring intervals $t$ from 2 to 64. The results revealed that intervals shorter than 8 led to near-ceiling performance, with participants achieving over 90% recall accuracy over all target images. Conversely, intervals exceeding 32 proved too challenging, with more than 30% of participants failing to identify any repeated images. Based on these findings, we selected the interval range of [8, 16, 24, 32] to ensure meaningful image-dependent variability in memory recall performance during the experiments. Each trial includes 100 unique target images, with foil images filling the sequence when targets are not shown, each target image and its repeat are shown only once to one participant. Sec. A.1 and Fig. A1 highlight differences in the experiment schematic between our study and previous works.
We select 2,500 images from ImageNet (Deng et al., 2009) as target images for the experiment, evenly distributed across five categories: dogs, cars, flowers, jellyfish, and teapots. A total of 3,263 participants from diverse demographics are recruited on Amazon Mechanical Turk (MTurk) (Turk, 2012). To ensure data quality, participants are disqualified if they incorrectly identify more than three foil images as target images in a row or fail to respond within 60 seconds. After filtering the data from disqualified participants, we analyzed responses from the remaining 1,000 qualified participants, resulting in 2,000 trials and a total of 171,568 key presses, with an average of 86 key presses per trial. The average false positive rate of the key presses is 14.67%. All experiments are conducted with informed consent and follow protocols approved by our institution’s Institutional Review Board.
@@ 3.2 Intra-Class Memorability Score (ICMscore)
To assess intra-class memorability, we introduce the Intra-Class Memorability Score (ICMscore), which measures the memorability of an image independently of its categorical information. Previous research typically computes memorability scores using hit rates and miss rates of memory recall (Bainbridge & Rissman, 2018; Goetschalckx & Wagemans, 2019; Kramer et al., 2023). However, the interval $t$ between image repetitions significantly affects memorability (Khosla et al., 2015). As shown in Fig. A2 and Sec. A.2, memory accuracy declines with longer intervals due to memory decays and increasing cognitive loads over time (Wixted, 2004; Sikarwar & Zhang, 2024). Therefore, we incorporate the lengths of these time intervals into the ICMscore calculation. Next, we describe how we compute ICMscore for a target image $I$.
@@
Since human memory retention diminishes over longer intervals, a longer interval $t$ makes correct identification of seen images more challenging. Therefore, when a participant correctly identifies an image after a longer interval, we assign it a higher memorability score $s$ to reflect the increased difficulty associated with the larger $t$. For correct responses ($x_{i}=1$), the memorability score $s_{i}$ is computed using a linearly increasing function: $s_{i}=x_{i}\frac{t_{i}}{T}$, where $t_{i}$ is the interval between two presentations of a target image, and $T=32$ is the maximum interval. Conversely, as longer time intervals are more prone to incorrect identifications of repeated target images, errors during shorter intervals are more indicative of the image’s memorability. For incorrect responses ($x_{i}=-1$), shorter intervals incur a larger penalty, while longer intervals result in a smaller penalty. The memorability score for incorrect responses is given by $s_{i}=x_{i}\frac{T-t_{i}}{T}$.
The final ICMscore for each target image $I$ is calculated as the average of its memorability scores $s_{i}$ over all $n$ participants: ICMscore(I)$=\frac{\sum_{i=1}^{n}s_{i}}{n}$. We also analyzed ICMscore without the temporal penalty $t_{i}$ (Sec. A.6). Results indicate that without this penalty, ICMscore fails to capture intra-class memorability, eliminating its impact on downstream computer vision tasks.
@@ 3.3 Dataset Analysis
Given the target images and their corresponding ICMscores, we now analyze how visual features of target images influence ICMscores. To investigate this relationship, we use a ResNet (He et al., 2015) model pre-trained on ImageNet (Deng et al., 2009) for object recognition tasks to extract feature representations for each image. We define the mean of all image features belonging to the same category as the centroid. Next, we compute the Pearson correlation (Sedgwick, 2012) between the L2 distances from its centroid to individual target image features and their ICMscores.
Figure 3: Correlation analysis between the feature distances of individual images to their categorical centroids and their corresponding ICMscore for the car and flower categories. Each “Car” and “Flower” image in ICMD is represented as a scatter point, with its corresponding ICMscore shown on the y-axis and the L2 distance from their centroids on the x-axis. The dashed lines, derived from linear regression, indicate a positive correlation between feature distance and CIMscore. As the L2 distance from the centroid increases, the ICMscore generally increases. The Pearson correlation coefficients are $r=0.46$ for “Car” and $r=0.42$ for “Flower” ($p<0.05$). See Sec. A.3 for the correlation analysis on other object categories.
Our results reveal a positive correlation between an image’s distance from its categorical centroid and its ICMscore across all object classes, as shown in Fig. 3 for the car and flower classes and Fig. A3 for the other object classes. The further away the features of an image are from its categorical centroid, the more memorable the image becomes. For example, a visually striking flower with distinct colors or intricate petal patterns stands out from typical flowers, making it more likely to be remembered (Fig. 1). Here, we explored the connections between memorability and feature distance within categories. Insights gained from this intra-class feature analysis motivate us to design the experiments that investigate the role of memorability in the following computer vision tasks in Sec. 4.